# Mitigating the Modality Gap: Few-Shot Out-of-Distribution Detection with Multi-modal Prototypes and Image Bias Estimation

### v1

@@ 4.1 RQ1: Multi-modal Prototypes
CLIP has been widely applied in different areas [58, 13]. However, a recent study [39] demonstrates that representations of images and texts are clearly separated, creating what is known as the modality gap. It remains unclear, however, whether this modality gap positively or negatively impacts OOD detection.
We hypothesize that the modality gap negatively impacts OOD detection performance by causing misalignment between image and text embeddings, and that this misalignment can lead to increased false positives, as OOD images may exhibit high similarity to ID text prototypes due to either semantic similarity or small spatial distance.
To test our hypothesis, we demonstrate that incorporating multi-modal (image and text) prototypes can mitigate the negative impact of the modality gap and reduce false positives. We obtain ID text prototypes $\{P_{t,c}\}_{c\in[C]}$ as in MCM [45]. For ID image prototypes $\{P_{i,c}\}_{c\in[C]}$, we collect ID base images $\mathcal{X}_{\textit{base}}=\{\mathbf{x}_{i}\}_{i\in[N_{\textit{base}}]}$ with corresponding labels $\{y_{\textit{base},i}\}_{i\in[N_{\textit{base}}]}$, where $N_{\textit{base}}$ is the number of base images. ID image prototypes are then calculated by averaging the image embeddings for each class,
$\small P_{i,c}=\frac{\sum_{i\in[N_{\textit{base}}]}\mathbb{I}(y_{\textit{base},i}=c)f_{\textit{image}}(\mathbf{x}_{i})}{\sum_{j\in[N_{\textit{base}}]}\mathbb{I}(y_{\textit{base},j}=c)},\forall c\in[C]\,,$ (3)
where $\mathbb{I}(\cdot)$ is the indicator function and $f_{\textit{image}}(\cdot)$ is CLIP’s image encoder. With these image prototypes, we extend $S_{\textit{MCM}}$ (Eq. 1[45] to a new OOD score, termed the Multi-Modal Prototypes score, $S_{\textit{MMP}}$, defined as,
$\small S_{\textit{MMP}}=\frac{S_{\textit{MCM}}(I_{t},\{P_{i,c}\}_{c\in[C]})+S_{\textit{MCM}}(I_{t},\{P_{t,c}\}_{c\in[C]})}{2}\,.$ (4)
@@
Table 1: Comparison of OOD detection results on ImageNet-100 (ID) $\rightarrow$ 4 OOD datasets (OOD). We use CLIP-B/16 for $S_{\textit{MCM}}$ and $S_{\textit{MMP}}$. FPR95 represents the false positive rate of OOD images when the true positive rate of ID images is at 95%. AUROC is the area under the receiver operating characteristic curve. Best in bold. Full results are in the Appendix.
Empirical evidence. We begin by empirically validating the effectiveness of image prototypes and $S_{\textit{MMP}}$. We use ImageNet-100 [7] as the ID dataset, with four OOD datasets: iNaturalist [59], SUN [63], Places [70], and Texture [6]. With all training images of ImageNet-100 as the base set, we observe in Tab. 1 that, on average, $S_{\textit{MMP}}$ outperforms $S_{\textit{MCM}}$ in both FPR95 and AUROC, demonstrating the effectiveness of image prototypes and $S_{\textit{MMP}}$.
Theoretical evidence. To further demonstrate why image prototypes improve performance, we present a theoretical justification in Theorem 1. The theorem shows that $S_{\textit{MMP}}$ increases the expected score separation between ID and OOD samples compared to using text prototypes alone as in $S_{\textit{MCM}}$. The proof is presented in the Appendix.
While incorporating image prototypes improves performance both empirically and theoretically, inspired by recent research [46, 36], we further fine-tune CLIP in a few-shot tuning manner to mitigate the impact of the modality gap (RQ2). The details are presented in the following section.
@@ 4.2 RQ2: suPreMe
To alleviate the negative impact of the image-text modality gap [39] in CLIP (RQ2), we propose suPreMe, which is summarized in Fig. 2. suPreMe incorporates two novel modules, i.e., the biased prompts generation module (BPG) and the image-text consistency module (ITC), along with a new OOD score. BPG enhances image-text fusion using image domain-biased prompts, while ITC minimizes both intra- and inter-modal distances. suPreMe’s novel OOD score, $S_{\textit{GMP}}$, leverages uni- and cross-modal similarities for improving robustness and performance.
Biased Prompts Generation (BPG).
Drawing inspiration from CoCoOp [73], we employ a set of learnable contexts $\{V_{i}\}_{i\in[L]}$ and the mapped image embedding $m(I_{t})$ to generate the text prototypes, where $L$ is the length of learnable contexts and $m(\cdot)$ is a two-layer MLP.
To enhance image-text fusion and improve generalization, we introduce a third component, a Gaussian-based estimated image domain bias $\mathcal{N}(\mathbf{\mu},\Sigma)$, for generating the prompts and ID text prototypes. This bias models the distribution of the mapped ID image embeddings, enabling the model to better distinguish between ID and OOD samples during inference. After that, we generate image domain-biased prompts (IDBP) conditioned on these three components. We next introduce the details of estimating the bias and generating IDBP.
Estimating image domain bias. In each training iteration, we sample $\mathbf{b}$ from the Gaussian-based image domain bias $\mathcal{N}(\mathbf{\mu},\Sigma)$ as,
@@
where $\Sigma=\sigma\sigma^{\top}$ (based on the Cholesky decomposition) and $\mathcal{N}(0,\mathcal{I})$ represents the standard Gaussian distribution, given identity matrix $\mathcal{I}$. To align the domain bias with the distribution of training ID image embeddings, we employ the following loss,
$\small\ell_{\textit{bias}}=\|\mu-m(I_{t})\|_{1}+\|\mathbf{b}-m(I_{t})\|_{1}\,,$ (7)
where $\|\cdot\|_{1}$ is the $\ell_{1}$ distance.
Generating image domain-biased prompts (IDBP). We condition our IDBP on three components: the learnable contexts $\{V_{i}\}_{i\in[L]}$, mapped image embedding $m(I_{t})$, and a sample $\mathbf{b}$ from the estimated domain bias $\mathcal{N}(\mathbf{\mu},\Sigma)$ as,
$\small\textit{IDBP}_{c}=[V_{1}+m(I_{t})+\mathbf{b},\ldots,V_{L}+m(I_{t})+\mathbf{b},y_{c}]\,.$ (8)
The final ID text prototype $P_{t,c}$ for the $c$-th class is obtained as $P_{t,c}=f_{\textit{text}}(\textit{IDBP}_{c})$.
Image-Text Consistency (ITC).
While BPG reduces the gap by image-text fusion, image-text consistency (ITC) directly reduces the modality gap by aligning image and text embeddings through inter- and intra-modal distances at a fine-grained level. It consists of two mappings, i.e., the image-to-text mapping $f_{i2t}(\cdot)$ and text-to-image mapping $f_{t2i}(\cdot)$, which are used for mapping embeddings into another domain/modality.
The first component of ITC is the inter-modal loss. It is designed to ensure that mapped text/image embeddings ($f_{t2i}(P_{t,c}),\forall c\in[C]$ and $f_{i2t}(I_{t})$) remain close to their original counterparts, thus enabling effective alignment. It is designed to minimize the cross-entropy loss,
$\small\ell_{\textit{inter}}=$ $-\log\frac{\exp({\operatorname{cos}(I_{t},f_{t2i}(P_{t,y}))/\tau})}{\sum_{c\in[C]}\exp({\operatorname{cos}(I_{t},f_{t2i}(P_{t,c}))/\tau})}$ (9)
$-\log\frac{\exp({\operatorname{cos}(f_{i2t}(I_{t}),P_{t,y})/\tau})}{\sum_{c\in[C]}\exp({\operatorname{cos}(f_{i2t}(I_{t}),P_{t,c})/\tau})}\,,$
where $y$ is the ground truth label for the input image, $C$ is the number of classes, and $\tau$ is the temperature.
Next, to prevent information loss, we introduce intra-modal loss. Specifically, we reconstruct the original embedding by mapping the mapped embedding back to its initial modality. Then, we minimize the $\ell_{1}$ distance between the reconstructed and original embeddings as follows,
$\small\ell_{\textit{intra}}=\|I_{t}-f_{t2i}(f_{i2t}(I_{t}))\|_{1}\\ +\sum_{c\in[C]}\|P_{t,c}-f_{i2t}(f_{t2i}(P_{t,c}))\|_{1}\,.$ (10)
The inter- and intra-modal losses serve to reduce the modality gap between image and text, and to preserve all information during mapping. To avoid overfitting, instead of using deep MLPs, we employ linear transformations without bias as the mappings ($f_{i2t}(\cdot)$ and $f_{t2i}(\cdot)$).
Combining BPG with ITC, suPreMe mitigates the modality gap by enhancing image-text fusion (BPG) as well as reducing inter- and intra-modal distances (ITC).
Training, Inference, and $S_{\textit{GMP}}$.
Generalized multi-modal prototypes OOD score $S_{\textit{GMP}}$. Building on our empirical and theoretical findings in Sec. 4.1, with the image-to-text mapping $f_{i2t}(\cdot)$, we introduce the generalized multi-modal prototypes OOD score $S_{\textit{GMP}}$. $S_{\textit{GMP}}$ uses the average of the maximum similarities between the multi-modal input embedding (the image embedding $I_{t}$ and mapped image embedding $I_{t}^{{}^{\prime}}=f_{t2i}(I_{t})$) and multi-modal prototypes:
$S_{\textit{GMP}}(I_{t},I_{t}^{{}^{\prime}},\{P_{t,c}\}_{c\in[C]},\{P_{i,c}\}_{c\in[C]})$ (11)
$=$ $\frac{1}{4}(S_{\textit{MCM}}(I_{t},\{P_{t,c}\}_{c\in[C]})+S_{\textit{MCM}}(I_{t},\{P_{i,c}\}_{c\in[C]})$
$+$ $S_{\textit{MCM}}(I_{t}^{{}^{\prime}},\{P_{t,c}\}_{c\in[C]})+S_{\textit{MCM}}(I_{t}^{{}^{\prime}},\{P_{i,c}\}_{c\in[C]}))\,.$
Remark.
$S_{\textit{GMP}}$ extends $S_{\textit{MCM}}$ by incorporating both uni- and cross-modal similarities. $S_{\textit{GMP}}$ balances contributions from both modalities, thus mitigating the negative impact of the modality gap. This approach enhances the separation between ID and OOD samples, improving the OOD detection performance of VLM-based methods as shown in Fig. 5.
Methods Venue ImageNet-1k $\rightarrow$ OOD Datasets Average
@@
MCM [45] NeurIPS’22 30.91 94.61 37.59 92.57 44.69 89.77 57.77 86.11 42.74 90.77
CoOp* [73] IJCV 29.47 94.89 31.34 93.36 40.28 90.07 54.25 87.58 38.83 91.47
CoCoOp* [72] CVPR’22 30.74 94.73 31.18 93.15 38.75 90.63 53.84 87.92 38.63 91.61
NPOS [57] ICLR’23 16.58 96.19 43.77 90.44 45.27 89.44 46.12 88.80 37.93 91.22
LoCoOp [46] NeurIPS’23 16.05 96.86 23.44 95.07 32.87 91.98 42.28 90.19 28.66 93.52
CATEX [40] NeurIPS’23 10.18 97.88 33.87 92.83 41.43 90.48 33.17 92.73 29.66 93.48
CLIPN [61] CVPR’23 23.94 95.27 26.17 93.93 33.45 92.28 40.83 90.93 31.10 93.10
NegLabel [24] ICLR’24 1.91 99.49 20.53 95.49 35.59 91.64 43.56 90.22 25.40 94.21
LSN [48] ICLR’24 21.56 95.83 26.32 94.35 34.48 91.25 38.54 90.42 30.22 92.96
EOE [4] ICML’24 12.29 97.52 20.40 95.73 30.16 92.95 57.53 85.64 30.09 92.96
ID-Like (4-shots) [2] CVPR’24 8.98 98.19 42.03 91.64 44.00 90.57 25.27 94.32 30.07 93.68
NegPrompt [36] CVPR’24 6.32 98.73 22.89 95.55 27.60 93.34 35.21 91.60 23.01 94.81
suPreMe 8.27 98.29 19.40 95.84 26.69 93.56 26.77 94.45 20.28 95.54
@@
Table 2: OOD detection results with ID data of ImageNet-1k and four OOD datasets using CLIP (VIT-B/16). Best in Bold. “*” is cited from LSN [48]. $\dagger$” represents results from MCM [45]. $\ddagger$” represents results from NPOS [57].
@@
Table 3: OOD detection results with ID data of ImageNet-100 and four OOD datasets using CLIP (VIT-B/16). Best in Bold. $\dagger$” represents results from LSN [48].

### v2

@@ 4.1 RQ1: Multi-modal Prototypes
CLIP has been applied in different areas [56, 13]. However, a recent study [39] demonstrates that representations of images and texts are clearly separated, creating what is known as the modality gap. It remains unclear, however, whether this gap positively or negatively impacts OOD detection.
We hypothesize that the modality gap negatively impacts OOD detection performance and that it can lead to increased false positives, as OOD images may exhibit high similarity to ID text prototypes due to either semantic similarity or small spatial distance.
To test our hypothesis, we demonstrate that incorporating multi-modal (image and text) prototypes can mitigate the negative impact of the modality gap and reduce false positives. We obtain ID text prototypes $\{P_{\textit{txt},c}\}_{c=1}^{C}$ as in MCM [44]. For ID image prototypes $\{P_{\textit{img},c}\}_{c=1}^{C}$, we collect ID base images $\{\mathbf{x}_{i}\}_{i=1}^{N_{\textit{base}}}$ with corresponding labels $\{y_{i}\}_{i=1}^{N_{\textit{base}}}$, where $N_{\textit{base}}$ is the number of base images. ID image prototypes are then calculated by averaging the image embeddings for each class label $y_{c}$:
$\small P_{\textit{img},c}=\frac{\sum_{i=1}^{N_{\textit{base}}}\mathbb{I}(y_{i}=y_{c})f_{\textit{image}}(\mathbf{x}_{i})}{\sum_{j=1}^{N_{\textit{base}}}\mathbb{I}(y_{j}=y_{c})},$ (3)
where $\mathbb{I}(\cdot)$ is the indicator function and $f_{\textit{image}}(\cdot)$ is CLIP’s image encoder. With these image prototypes, we extend $S_{\textit{MCM}}$ (Eq. 1[44] to a new OOD score, which we term the Multi-Modal Prototypes score, $S_{\textit{MMP}}$, defined as
$\small S_{\textit{MMP}}=\frac{S_{\textit{MCM}}(I,\{P_{\textit{img},c}\}_{c=1}^{C})+S_{\textit{MCM}}(I,\{P_{\textit{txt},c}\}_{c=1}^{C})}{2}\,.$ (4)
@@
Table 1: Comparison of OOD detection results on ImageNet-100 (ID) $\rightarrow$ 4 OOD datasets (OOD) with CLIP-B/16. FPR95 represents the false positive rate of OOD images when the true positive rate of ID images is at 95%. AUROC is the area under the receiver operating characteristic curve. Best in bold.
Empirical evidence. We begin by empirically validating the effectiveness of image prototypes and $S_{\textit{MMP}}$. We use ImageNet-100 [6] as the ID dataset, with four OOD datasets: iNaturalist [57], SUN [61], Places [71], and Texture [5]. With all training images of ImageNet-100 used for obtaining image prototypes, we observe in Tab. 1 that, on average, $S_{\textit{MMP}}$ outperforms $S_{\textit{MCM}}$ in FPR95 and AUROC, demonstrating the effectiveness of image prototypes and $S_{\textit{MMP}}$.
Theoretical observation. To further understand why image prototypes improve performance, we present a theoretical analysis in Theorem 1 of the Appendix. The theorem shows that $S_{\textit{MMP}}$ increases the expected score separation between ID and OOD samples compared to using text prototypes alone (as in $S_{\textit{MCM}}$) under general assumptions on OOD and ID distributions used by previous works [12, 39].
While incorporating image prototypes improves performance , inspired by recent research [45, 36], we further fine-tune CLIP in a few-shot tuning manner to mitigate the impact of the modality gap (RQ2).
@@ 4.2 RQ2: suPreMe
To alleviate the negative impact of the image-text modality gap [39] in CLIP (RQ2), we propose suPreMe, which is summarized in Fig. 3. suPreMe incorporates two modules, i.e., the biased prompt generation module (BPG) and the image-text consistency module (ITC), along with a new OOD score. BPG enhances image-text fusion using image domain-biased prompts, while ITC minimizes both intra- and inter-modal distances. suPreMe’s novel OOD score, $S_{\textit{GMP}}$, exploits uni- and cross-modal similarities to improve robustness and performance.
Biased Prompt Generation (BPG).
Drawing inspiration from CoCoOp [74], we employ a set of learnable context tokens $\{V_{i}\in\mathcal{R}^{1\times N_{\text{lm}}}\}_{i=1}^{L}$ and the mapped image embedding $m(I)\in\mathcal{R}^{1\times N_{\text{lm}}}$ to generate the text prototypes, where $L$ is the length of learnable contexts, $m(\cdot)$ is a two-layer MLP, and $N_{\text{lm}}$ is the dimension of text encoder’s input. Similar to previous methods [73, 45], learnable context tokens replace the manually designed prompt templates, e.g., “a photo of a class ”, for generating text prototypes.
To improve generalization on unseen ID data and avoid overfitting on few-shot training data, we introduce a third component, a Gaussian-based estimated image domain bias $\mathcal{N}(\mathbf{\mu},\Sigma)$. This bias counteracts the modality gap by pre-biasing text prototypes toward ID images and augments the training data to better approximate the true distribution of mapped ID image embeddings. Using a Gaussian distribution is a reasonable a priori choice and is mathematically tractable. Our ablation study (Tab. 8) confirms that this choice is valid and is successful in reducing FPs. After that, we generate image domain-biased prompts (IDBP) conditioned on these three components.
Estimating image domain bias. In each training iteration, we sample $\mathbf{b}\in\mathcal{R}^{1\times N_{\text{lm}}}$ from $\mathcal{N}(\mathbf{\mu},\Sigma)$ as
@@
where $\Sigma=\sigma\sigma^{\top}$ (based on the Cholesky decomposition) and $\mathcal{N}(0,\mathcal{I})$ represents the standard Gaussian distribution, given identity matrix $\mathcal{I}$. To align the domain bias with the distribution of training ID image embeddings, we employ the following loss:
$\small\ell_{\textit{bias}}(I)=\|\mu-m(I)\|_{1}+\|\mathbf{b}-m(I)\|_{1}\,$, (6)
where $\|\cdot\|_{1}$ is the $\ell_{1}$ distance. In the first term of the loss (Eq. 6), $\mu$ is updated. In the second term, as $\mathbf{b}$ is sampled using $\mu$ and $\Sigma$, $\mu$ and $\Sigma$ will be updated during the few-shot training procedure. The first term, $\mu-m(I)||_{1}$, pushes the bias distribution’s mean $\mu$ toward the centroid of training image embeddings, while the second term, ${\mathbf{b}}-m(I)||_{1}$, ensures sampled bias vectors $\mathbf{b}$ are within the training image embedding space.
Methods Venue Shots ImageNet-1k $\rightarrow$ OOD Datasets Average
@@
MCM [44] NeurIPS’22 0 30.91 94.61 37.59 92.57 44.69 89.77 57.77 86.11 42.74 90.77
CLIPN [59] CVPR’23 0 23.94 95.27 26.17 93.93 33.45 92.28 40.83 90.93 31.10 93.10
EOE [3] ICML’24 0 12.29 97.52 20.40 95.73 30.16 92.95 57.53 85.64 30.09 92.96
CoOp* [74] IJCV 16 29.47 94.89 31.34 93.36 40.28 90.07 54.25 87.58 38.83 91.47
CoCoOp* [73] CVPR’22 16 30.74 94.73 31.18 93.15 38.75 90.63 53.84 87.92 38.63 91.61
NPOS [55] ICLR’23 - 16.58 96.19 43.77 90.44 45.27 89.44 46.12 88.80 37.93 91.22
LoCoOp [45] NeurIPS’23 16 16.05 96.86 23.44 95.07 32.87 91.98 42.28 90.19 28.66 93.52
CATEX [40] NeurIPS’23 - 10.18 97.88 33.87 92.83 41.43 90.48 33.17 92.73 29.66 93.48
LSN [47] ICLR’24 - 21.56 95.83 26.32 94.35 34.48 91.25 38.54 90.42 30.22 92.96
ID-Like [2] CVPR’24 4 8.98 98.19 42.03 91.64 44.00 90.57 25.27 94.32 30.07 93.68
NegPrompt [36] CVPR’24 16 6.32 98.73 22.89 95.55 27.60 93.34 35.21 91.60 23.01 94.81
SCT [66] NeurIPS’24 16 13.94 95.86 20.55 95.33 29.86 92.24 41.51 89.06 26 47 93.37
Local-Prompt [67] ICLR’25 16 8.63 98.07 23.23 95.12 31.74 92.42 34.50 92.29 24.52 94.48
suPreMe 16 5.91 98.73 15.85 96.08 26.18 93.19 34.88 91.93 20.70 95.38
suPreMe 16 8.27 98.29 19.40 95.84 26.69 93.56 26.77 94.45 20.28 95.54
@@
Table 2: OOD detection results with ID data of ImageNet-1k and four OOD datasets using CLIP (VIT-B/16). Best in bold. Second best is underlined. “*” is cited from LSN [47]. $\dagger$” represents results from MCM [44]. $\ddagger$” represents results from NPOS [55]. suPreMe employs 16 images per class for training and the whole training split for calculating ID image prototypes.
Generating image domain-biased prompts (IDBP). We condition IDBPs on three components: the learnable contexts $\{V_{i}\in\mathcal{R}^{1\times N_{\text{lm}}}\}_{i=1}^{L}$, mapped image embedding $m(I)\in\mathcal{R}^{1\times N_{\text{lm}}}$, and a sample $\mathbf{b}\in\mathcal{R}^{1\times N_{\text{lm}}}$ from $\mathcal{N}(\mathbf{\mu},\Sigma)$ as
$\small\textit{IDBP}_{c}=[V_{1}+m(I)+\mathbf{b},\ldots,V_{L}+m(I)+\mathbf{b},y_{c}]\,.$ (7)
The final ID text prototype $P_{\textit{txt},c}$ for the $c$-th class is obtained as $P_{\textit{txt},c}=f_{\textit{text}}(\textit{IDBP}_{c})$.
Image-Text Consistency (ITC).
While BPG reduces the gap by image-text fusion, image-text consistency (ITC) directly reduces the modality gap by aligning image and text embeddings through inter- and intra-modal distances at a fine-grained level. It consists of two mappings, i.e., the image-to-text mapping $f_{\textit{img-txt}}(\cdot)$ and text-to-image mapping $f_{\textit{txt-img}}(\cdot)$, which are used for mapping embeddings into the other modality.
The first component of ITC is the inter-modal loss. It is designed to ensure that mapped text/image embeddings ($\{f_{\textit{txt-img}}(P_{\textit{txt},c})\}_{c=1}^{C}$ and $f_{\textit{img-txt}}(I)$) remain close to their original counterparts, thus enabling effective alignment. It is designed to minimize the cross-entropy loss:
$\ell_{\textit{inter}}(I)=$ $-\log\frac{\exp({\cos(I,f_{\textit{txt-img}}(P_{\textit{txt},c}))/\tau})}{\sum_{j=1}^{C}\exp({\cos(I,f_{\textit{txt-img}}(P_{\textit{txt},j}))/\tau})}$ (8)
$-\log\frac{\exp({\cos(f_{\textit{img-txt}}(I),P_{\textit{txt},c})/\tau})}{\sum_{j=1}^{C}\exp({\cos(f_{\textit{img-txt}}(I),P_{\textit{txt},j})/\tau})}\,$,
where $c$ is the index of the ground truth label class for the input image and $\tau$ is the temperature.
Next, to prevent information loss, we introduce intra-modal loss. Specifically, we reconstruct the original embedding by mapping the mapped embedding back to its initial modality. Then, we minimize the $\ell_{1}$ distance between the reconstructed and original embeddings as follows:
$\ell_{\textit{intra}}(I)=$ $\|I-f_{\textit{txt-img}}(f_{\textit{img-txt}}(I))\|_{1}$ (9)
$+\sum_{j=1}^{C}\|P_{\textit{txt},j}-f_{\textit{img-txt}}(f_{\textit{txt-img}}(P_{\textit{txt},j}))\|_{1}\,.$
The inter- and intra-modal losses serve to reduce the modality gap between image and text, and to preserve all information during mapping.
Training, Inference, and $\boldsymbol{S_{\textit{GMP}}}$.
Generalized multi-modal prototypes OOD score $\boldsymbol{S_{\textit{GMP}}}$. Building on our empirical and theoretical findings in Sec. 4.1, with the image-to-text mapping $f_{\textit{img-txt}}$, we introduce the generalized multi-modal prototypes OOD score $S_{\textit{GMP}}$. $S_{\textit{GMP}}$ uses the average of the maximum similarities between the multi-modal input embedding ($I$ and $I^{\prime}=f_{\textit{txt-img}}(I)$) and multi-modal prototypes:
$S_{\textit{GMP}}$ $(I,I^{\prime},\{P_{\textit{txt},c}\}_{c=1}^{C},\{P_{\textit{img},c}\}_{c=1}^{C})=$
$\big(S_{\textit{MCM}}(I,\{P_{\textit{txt},c}\}_{c=1}^{C})+S_{\textit{MCM}}(I,\{P_{\textit{img},c}\}_{c=1}^{C})$ (10)
$+$ $S_{\textit{MCM}}(I^{\prime},\{P_{\textit{txt},c}\}_{c=1}^{C})+S_{\textit{MCM}}(I^{\prime},\{P_{\textit{img},c}\}_{c=1}^{C})\big)/4\,.$
Remark.
$S_{\textit{GMP}}$ extends $S_{\textit{MCM}}$ by incorporating both uni- and cross-modal similarities. In different scenarios, the importance of intra- and inter-modal similarity is different. However, we average the contributions from the image and text modalities as a starting point to ensure the generalization ability of $S_{\textit{GMP}}$ and reduce the gap. This approach enhances the separation between ID and OOD samples, improving the OOD detection performance of VLM-based methods, as shown in Fig. 5.
@@
Table 3: OOD detection results with ID data of ImageNet-100 and four OOD datasets using CLIP (VIT-B/16). Best in bold. Second best is underlined.